### v4

# Advances in Embodied Navigation Using Large Language Models: A Survey

Abstract
In recent years, the rapid advancement of Large Language Models (LLMs) such as the Generative Pre-trained Transformer (GPT) has attracted increasing attention due to their potential in a variety of practical applications. The application of LLMs with Embodied Intelligence has emerged as a significant area of focus. Among the myriad applications of LLMs, navigation tasks are particularly noteworthy because they demand a deep understanding of the environment and quick, accurate decision-making. LLMs can augment embodied intelligence systems with sophisticated environmental perception and decision-making support, leveraging their robust language and image-processing capabilities. This article offers an exhaustive summary of the symbiosis between LLMs and embodied intelligence with a focus on navigation. It reviews state-of-the-art models, research methodologies, and assesses the advantages and disadvantages of existing embodied navigation models and datasets. Finally, the article elucidates the role of LLMs in embodied intelligence, based on current research, and forecasts future directions in the field. A comprehensive list of studies in this survey is available at https://github.com/Rongtao-Xu/Awesome-LLM-EN.
keywords:
Large Language Models, Embodied Intelligence, Navigation.
journal: Engineering Applications of Artificial Intelligence
1 Introduction
The development of Large Language Models (LLMs) in the field of navigation is rapidly evolving and holds significant potential for advancing natural language processing and machine learning. Notably, LLMs have already achieved remarkable success in Few-Shot Planning, enabling effective planning and decision-making for new tasks with minimal or no sample data. However, alongside these achievements, numerous technical and theoretical challenges persist. These challenges include the integration of text, images, and other sensor data simultaneously, the reduction of latency for real-time applications, and the enhancement of training efficiency without sacrificing performance.
To address these challenges, researchers have employed a diverse array of methods such as machine learning [1], deep learning [2] and evolutionary algorithms [3] [4], as well as the addition of other vision-language models to enhance the multimodal data fusion capabilities of LLMs. These methodologies aim to develop navigation agents capable of learning from experience and continually improving their performance. In this paper, we review existing models that have employed LLM-based agents for navigation tasks. These studies have demonstrated that LLM-based agents perform excellently in path planning, environmental understanding, and dynamic adjustment tasks. The success of these studies underscores the utility of LLM-based agents as valuable tools for navigation.
This paper primarily explores the usage of LLMs in navigation models. By delving into these technologies, we provide comprehensive knowledge and resources for researchers and practitioners. Compared to other related reviews, this paper focuses specifically on the application of large language models in the field of navigation, analyzing the strengths and weaknesses of existing LLM-based navigation models, and comparing them with non-LLM-based Vision-and-Language Navigation (VLN) models [5] [6] [7]. Additionally, we review common datasets in the navigation field, compare these datasets, and highlight the current issues with them. Our contributions are as follows:
  1. 1.
    Comprehensive Review and Comparative Analysis: The article provides a comprehensive review of the application of LLMs in the navigation field, covering multiple LLM-based navigation models and comparing them with non-LLM-based models. This integrated analysis helps readers fully understand current research developments and highlights the advantages and disadvantages of LLMs in navigation.
  2. 2.
    Dataset Analysis: The article thoroughly analyzes commonly used datasets in the navigation field, discussing their applicability, issues, and limitations. This detailed analysis guides researchers in selecting appropriate datasets and offers practical advice for their application, enhancing the article’s practical relevance.
  3. 3.
    Challenges and Future Directions: The article not only showcases current research advancements but also delves into the challenges and issues faced by existing methods in practical applications. By identifying these challenges, the article points out directions for future research and highlights innovative application scenarios and technological trends of LLMs in the navigation field.
2 Background
2.1 Large Language Models
The development of LLMs represents a significant milestone in the domains of Natural Language Processing (NLP) [8] and Machine Learning. LLMs are machine learning models specifically designed to process and generate natural language [9]. They are typically trained on extensive text data to capture the grammatical structures, inter-word relationships, and contextual nuances of language. These models predominantly utilize the Transformer architecture, which is distinguished by its self-attention mechanism that effectively addresses long-range dependencies. By employing pre-training on broad corpora to learn fundamental language patterns followed by fine-tuning on specialized datasets for specific tasks, LLMs are adapted for diverse applications. Additionally, these models are capable of not only understanding complex textual information but also generating coherent and fluent text outputs, demonstrating their high efficacy in natural language generation, translation, summarization, and question answering. Key technical features of LLMs include:
  1. 1.
    Word Embedding: Models process text by converting words into numerical vectors, known as word embeddings, which encapsulate the semantic and syntactic properties of vocabulary, laying the groundwork for further language processing.
  2. 2.
    Layers and Heads: The Transformer architecture comprises multiple layers, each containing several ’heads’ (Attention Heads) that process different types of information in parallel, enhancing the model’s capacity to handle information.
  3. 3.
    Sequence Processing: Unlike Recurrent Neural Networks, Transformers can process entire input sequences in parallel, significantly improving efficiency and speed.
Tracing the evolution of these models requires examining the early stages of NLP and machine learning, where methods such as Bag-of-Words (BoW) [10] were prevalent for text representation but limited in capturing word order and contextual nuances.
Before the mainstream adoption of deep learning, simpler algorithms like N-grams and decision trees held sway in a variety of language-processing tasks. N-grams served as foundational elements for text and language data analysis by breaking down text into sequences of N contiguous words or characters. Although efficient and straightforward, N-grams suffer from limitations such as failing to capture long-distance dependencies and requiring a large parameter space for substantial N-values.
Decision trees have been employed for classification and regression tasks. These trees recursively partition the dataset, where each internal node denotes a feature, each branch symbolizes a decision rule, and each leaf node represents the final output. While decision trees offer ease of interpretation and handle irrelevant features effectively, they are susceptible to overfitting and struggle with non-linear, high-dimensional data.
The advent of word embedding models around 2013, notably Word2Vec and GloVe, signaled a transformative shift in NLP. These models allowed for words to be mapped into a continuous vector space, thereby capturing intricate word relationships with remarkable effectiveness. With the advent of word embedding techniques , the focus shifted toward incorporating these pre-trained word vectors into sequence models like Recurrent Neural Networks (RNNs) and LSTMs. RNNs are neural architectures specifically designed to handle sequence data by utilizing the temporal context from prior elements in a sequence for processing subsequent inputs. Nevertheless, during their training, RNNs are plagued by the vanishing and exploding gradient problems, which hinder their capacity to capture long-term dependencies and contribute to training instability.Both models have their merits and drawbacks, but they are now primarily used as baseline models or for feature engineering in the wake of more complex models enabled by deep learning.
To mitigate these drawbacks, specialized RNN variants such as Gated Recurrent Units (GRUs) [11] have gained prominence in sequence data analysis, particularly within NLP. LSTMs, equipped with gating mechanisms, effectively circumvent the vanishing gradient issue but at the cost of added computational complexity and the potential for exploding gradients. The increased number of parameters also increases the risk of overfitting and extends training time.
Building on word embeddings and sequence models, a new paradigm emerged with the development of end-to-end learning architectures, most notably the Transformer model. Introduced by Google in 2017, the Transformer architecture eschews recurrent and convolutional layers, opting instead for a Self-Attention mechanism to capture sequence dependencies [12]. This design enables parallel data processing, resulting in a significant speed-up over RNNs and LSTMs. The Transformer’s scalability also benefits from its layered architecture, facilitating the addition of layers or parameters to improve performance.
Following the Transformer, Google’s Vision Transformer (ViT) [13] in 2020 and OpenAI’s CLIP [14] in 2021 marked the model’s extension from NLP to visual processing tasks. Both architectures employ Transformer-based text encoders and demonstrate the model’s versatility across domains.Among the various Transformer-based architectures, Bidirectional Encoder Representations from Transformers (BERT) [15], introduced by Google in 2018, garnered attention for its pre-training approach across diverse NLP tasks. Employing bidirectional encoders, BERT captures both preceding and succeeding contextual information for each word. While effective, BERT’s complexity results in high computational requirements. OpenAI’s introduction of the GPT series [16] [17] [18] has been another milestone, particularly for text generation tasks. These models also exhibit ”zero-shot learning” capabilities, enabling them to perform certain tasks without task-specific training. However, for tasks requiring higher accuracy, fine-tuning is generally necessary.
GPT models have been deployed in a wide array of applications in NLP and artificial intelligence, such as chatbots, automated writing, and question-answering systems, and have opened new avenues in multi-modal learning and code generation.
In summary, LLMs are increasingly shaping various domains, and their ongoing development continues to reveal new potential. LLMs have been widely applied in multiple areas including, but not limited to, natural language processing tasks [9], autonomous driving [5], robotics task planning [19], and knowledge-intensive tasks [20]. Particularly in the fields of autonomous driving and robotics task planning, LLMs help develop precise and efficient action plans by understanding natural language instructions and environmental information [21]. For instance, in the context of self-driving vehicles or mobile robots, LLMs can assist the system in understanding descriptions of destinations and generating route guides to reach specified locations. Moreover, these models can handle user queries, such as information requests about nearby restaurants or specific places, thereby enhancing the user experience of interactive navigation systems. In embodied systems, by integrating LLMs, devices can not only interact more intelligently with users but also adapt and make efficient decisions in a broader range of environments, showcasing the integration and innovation of technology.The widespread deployment and diverse applications of these models not only underscore their immediate utility but also open exciting possibilities for future research.
2.2 Embodied Intelligence
Embodied Intelligence is an emerging field focused on understanding and developing intelligent agents that closely interact with their environment. This field posits that intelligence fundamentally emerges from an agent’s interaction with its surroundings, rather than being an inherent characteristic of isolated, abstract computations. Drawing on various disciplines such as neuroscience, psychology, robotics, and artificial intelligence, Embodied Intelligence seeks to create innovative models and algorithms for simulating intelligent behavior [22].
In the recent developments within the field of Embodied Intelligence, the integration of LLMs with intelligent agents has become a hot topic. This approach leverages the extensive knowledge and excellent NLP capabilities of large language models, translating human instructions into formats interpretable by physically embodied agents. Although large language models excel in natural language processing, they typically lack direct interaction with the physical world, as their training data is predominantly textual. To enhance the efficiency and generalization ability of these agents, many researchers are no longer limiting themselves to using only text as input. Instead, they are integrating images, point clouds [23] [24], and even voice prompts into their projects. Including using finetuned large language models and training with methods such as reinforcement learning [25] [26], these ways undoubtedly align with the concept of Embodied Intelligence and marks a progression towards the integration of multimodal algorithms [27]. Multimodality refers to the integration and utilization of multiple sensory modalities in perception, communication, and cognition. By leveraging multimodal information, such as visual, auditory, and tactile cues, embodied agents can gather a more comprehensive understanding of their environment . This multimodal perception enables them to make sense of complex spatial information, recognize objects and events, and navigate effectively in real-world scenarios [28].
LLMs such as GPT and other multimodal models have demonstrated significant capabilities in processing both images and language, integrating natural language processing and computer vision technologies. These models are trained on vast datasets of images and corresponding textual descriptions, learning to associate visual content with natural language, which allows them to excel in various tasks like image classification and caption generation. Particularly in embodied intelligence systems, such as smart homes, robots, or autonomous vehicles, the image and language processing abilities of LLMs can significantly enhance user interaction. The systems analyze images and leverage language models to understand user needs, enabling smart responses. Some multimodal LLMs process images by feeding them into various image encoders, such as the Vision Transformer [13], and then combine them with relevant context into a sequence of tokens, which are then input into the corresponding LLM for image recognition and processing. Additionally, models like DALL-E allow users to generate new images or edit existing ones by inputting descriptive text [29]. For example, a user might request ”a cat in a spacesuit on the moon,” and the model can generate an image based on this description. LLMs also provide decision support in scenarios requiring real-time information processing and adapt to specific environments or needs through interaction with users. This combination of advanced image processing technology and sophisticated language understanding enables LLMs to be extensively applied in enhancing embodied intelligence systems, offering unprecedented levels of interaction and automation.
Besides, In the realm of Embodied Intelligence, the concept of an intelligent agent is a pivotal element. Control over these agents is divided into high-level and low-level aspects. High-level controls involve task scheduling and strategy development, incorporating techniques like reinforcement learning [30], deep learning [31] [32] [33] [34] [35], and emerging methodologies based on expansive language models. In contrast, low-level control relates to direct commands over the agent’s operational functions, such as position, speed, and force. An agent’s capabilities depend on its design and low-level control parameters [36], while high-level controls orchestrate the execution and approach to task completion [37].
These methodologies have led to a range of distinct applications, including terrain recognition [38], predicting machinery lifespan [39], and emulating gaze mechanisms [40]. These developments not only reflect the dynamic nature of the field but also signify the tangible impact of these controlled embodied agents across various domains. The advancement of intelligent agents hinges on the harmonious integration of high-level and low-level controls. Utilizing superior algorithms and control techniques is essential for developing new agents characterized by greater robustness and broader generalization capabilities. In addition to research on single-agent systems, studies on multi-agent systems focus independently on multi-agent cooperation issues [41] [42] [43]. Their research efforts are instrumental in expanding the range of tasks achievable by agents, enhancing work efficiency, and improving the real-world applicability and generalization of Embodied Agents [44].
To encapsulate, the field of Embodied Intelligence is navigating a transformative journey: starting from reinforcement learning, shifting towards the utilization of LLMs, and currently progressing into the realm of multimodal integration [45]. As a subset of Artificial Intelligence, the advancement of Embodied Intelligence is intrinsically tied to the practical deployment of diverse algorithms within the physical world.
3 LLMs in Embodied Agents
3.1 LLMs for Grounded Language Understanding
Grounded Language Understanding aims to reconcile the abstract symbols processed by language models with concrete entities, actions, or states in the physical or simulated world. This is pivotal for applications requiring real-world interaction, such as robotic control, natural language interfaces, or advanced research in Embodied Intelligence.
Within this context, LLMs like GPT and BERT can integrate with sensors, databases, or simulated environments to generate and interpret language applicable to real-world scenarios. In a robotic setting, for instance, an LLM could interpret sensor data, process natural language directives, and issue control signals to the robot, thereby bridging high-level language and low-level actions.
LLMs’ application in grounded language understanding is an evolving research area, intersecting with disciplines such as robotics, computer vision, and human-computer interaction. Techniques like Few-Shot Learning and Transfer Learning are commonly used to adapt these pre-trained models to specialized grounding tasks with minimal additional training.
Although LLMs have advanced significantly in NLP, they primarily excel in text generation or classification and seldom engage with tangible entities or real-world situations. Their limitations include a lack of foundational knowledge, impairing their ability to process interrelated concepts and function in interactive settings (Mahowald et al. [46]). To mitigate these issues, researchers have employed various strategies, such as 1) fine-tuning pre-trained models, 2) implementing reinforcement learning algorithms for decision-making in complex environments, and 3) devising specialized architectures for multimodal learning.
Despite their immense potential and practical applications, LLMs face numerous technical and theoretical challenges in grounded language understanding. These include the integration of text, images, and sensor data, latency reduction for real-time applications, and maintaining training efficiency without sacrificing performance. With ongoing research and emerging technologies, significant advancements in this domain are anticipated.
3.2 LLMs for Few-Shot Planning
LLMs have achieved notable success in NLP and are increasingly being applied to Few-Shot Planning, a machine learning technique designed for effective planning and decision-making in new tasks with minimal sample data. LLMs excel in zero-shot and few-shot learning, allowing for efficient generalization across tasks without the need for task-specific training or limited examples. Employing LLMs for Few-Shot Planning often involves processing natural language queries, enabling the models to generate actionable plans or steps, thus proving invaluable in zero-shot navigation tasks.
As discussed in research by Zhou et al. in ESC [47] and Song et al. in LLM-Planner [48], the effectiveness of navigation tasks is closely related to semantic scene understanding, which relies heavily on Few-Shot Planning. Planners incorporating LLMs are increasingly vital for enabling embodied agents to execute complex tasks in visually rich environments based on natural language directives. The performance of such planners has a direct bearing on the task’s overall success. The role of LLMs as planners in embodied tasks is expanding, thanks in part to the emergence of methodologies like LLM-Planner [48], CoT (Chain-of-Thought)[49], and LLM-DP[50]. These methods complement the already respectable performance of traditional symbolic planners like the Fast-Forward planner [51] and BFS(f) planner [52] in current research.
3.3 Embodied Navigation
In the field of embodied navigation, state-of-the-art models enhance their performance and practicality through various methods. Firstly, these models achieve multimodal integration [45], combining visual, linguistic, and auditory data, using LLMs to process complex language instructions in conjunction with visual systems to accurately interpret and respond to the environment [53]. Secondly, they enhance understanding of environmental context and decision-making through deep learning algorithms [2], leveraging the advanced language comprehension capabilities of LLMs to better grasp task requirements and navigate environmental obstacles [54]. Additionally, models with long-term memory mechanisms can remember past navigation experiences and environmental features, aiding effective path planning in complex or changing environments [54] [55]. These models also possess the ability to interact effectively with users, processing natural language inputs to communicate in a more natural and human-like manner. Lastly, the adoption of reinforcement learning methods allows models to self-optimize and learn through continuous interaction with the environment, continually adjusting behaviors to adapt to new tasks or environmental changes. By integrating these technologies, state-of-the-art embodied navigation models not only improve navigation efficiency but also enhance their ability to cope with complex environments, achieving highly intelligent and adaptable navigation capabilities [6].
3.3.1 LLM-based Model
In recent years, a rapidly expanding array of works has emerged in the field of navigation, each characterized by its unique emphasis and methodologies for performance assessment. Generally, these works employ approaches that fall into two categories, as illustrated in Fig.2 and Fig.3. The first category uses LLMs as planners, directly generating actions and leveraging exploration policies to guide agents. The second category employs LLMs to scrutinize incoming visual or textual data to isolate goal-relevant information, based on which exploration policies subsequently produce suitable actions for agent navigation. An architectural diagram delineating these approaches is presented below. This section endeavors to examine the nuanced features and distinctions among these works, furnishing a comprehensive temporal map of the works under discussion in this chapter, as depicted in Fig.1. In doing so, we aim to offer an exhaustive understanding that will prove invaluable for both scholars and professionals in this field.
Figure 1: This presentation exhibit a temporal map depicting the works of embodied navigation from 2022 to 2023, and we selected 5 typical works to showcase their corresponding framework diagrams. The map illustrates the evolution of major works, offering valuable insights into the advancement of Embodied Agents.
Table 1: Classification of the works described in this section.
Categories Works
Planning CLIP-Nav NavGPT VELMA $A^{2}$av MiC SayNav
Semantic Understanding LM-Nav L3MVN ESC SQA3D
1)LM-Nav [56] is a navigation architecture designed for robots that leverages pre-existing models specialized in language, vision, and action to enable sophisticated interactions with robots via natural language commands. Remarkably, the system eliminates the need for costly supervision and fine-tuning, relying solely on pre-trained models for navigation, image-langu-age correlation, and language modeling. The architecture of LM-Nav consists of three integrated, pre-trained models to ensure precise instruction execution in complex, real-world scenarios. Specifically, the system employs GPT-3 as the LLM, tasked with decoding verbal instructions into a series of textual landmarks. Concurrently, CLIP serves as the VLM, anchoring these textual landmarks to a topological map. Lastly, the Vision-Action Model (VAM) is a self-supervised robotic control model, responsible for utilizing visual data and executing physical actions based on plans synthesized by the LLM and VLM.
Implemented on a real-world mobile robot, LM-Nav has been shown to accomplish long-horizon navigation in intricate, outdoor settings solely based on natural language directives. A salient feature is the lack of model fine-tuning; all three component models are trained on expansive datasets with self-supervis-ed learning objectives and are deployed as-is. The system has demonstrated its ability to interpret and execute natural language instructions across significant distances in complex, suburban terrain while also offering disambiguation in path selection through detailed commands. Such performance metrics underscore LM-Nav’s robust generalization capabilities and its proficiency in navigating complicated outdoor environments.
2)CLIP-NAV [57] introduces an innovative ”zero-shot” navigation framework aimed at solving coarse-grained instruction-following tasks. The architecture is structured to dissect the guidance language into critical keyphrases, visually anchor them, and leverage the resulting grounding scores to direct the CLIP-Nav sequence-to-sequence model in predicting the agent’s subsequent actions. Initially, a keyphrase extraction module isolates salient terms from the given instructions. Following this, a visual grounding module anchors these keyphrases within the environment, thereby generating a set of grounding scores. These scores serve as the basis for the sequence-to-sequence model in CLIP-Nav, which computes the next set of actions based on the agent’s current state and grounding scores.
To augment the efficacy of CLIP-Nav, the model incorporates a backtracking mechanism. This allows the agent to retrace its steps, facilitating revisions to prior decisions. Such a feature is particularly beneficial in the context of coarse-grained instruction-following tasks, wherein corrective backtracking may be essential.
Evaluation metrics confirm the robustness of CLIP-Nav, establishing a zero-shot baseline on the REVERIE task. Remarkably, CLIP-Nav exceeds the performance of the unseen supervised baseline, even without dataset-specific fine-tuning, in both success rate and success weighted by path length. Additionally, the paper introduces a new performance metric, termed Relative Change in Success (RCS), to quantitatively assess generalizability in vision and language navigation tasks. The RCS metric substantiates the superior performance of CLIP-based methodologies over conventional supervised approaches.
Figure 2: The first type employs LLMs as planners that directly generate actions, thereby leveraging exploration policies to control agents. LLMs, by acquiring information from text and other visual models processed through images, perform planning (using a dialogue format as an example here) and hand the actions over to exploration algorithms and agents for navigation to complete tasks.
3)SQA3D [58] (Situated Question Answering in 3D Scenes) introduces a task formulated to assess the scene comprehension aptitude of embodied agents. The task necessitates that the agent garner an exhaustive understanding of its orientation within a 3D environment, guided by a text-based description, and subsequently generate precise answers to questions pertaining to that understanding.
The principal objective of SQA3D is to gauge the capacity of embodied agents to engage in logical reasoning about their immediate environment and generate answers based on such reasoning. Unlike most existing tasks, which presume that observations are made from a third-person viewpoint, SQA3D uniquely demands that agents construct and reason from an egocentric perspective of the scene.
SQA3D is underpinned by an extensive dataset derived from 650 scenes in ScanNet, comprising 6.8k distinct scenarios, 20.4k textual descriptions, and 33.4k multifaceted reasoning questions. The scope and diversity of the dataset enable the exploration of complex questions that demand a nuanced understanding and facilitate the aggregation of a substantial data corpus.
To address SQA3D’s inherent challenges, researchers employed transformer-based vision-language models such as Scan-QA, which generates answers by associating 3D scans with questions. In addition, auxiliary tasks have been considered during the training phase to amplify model performance. Specialized models attuned to various 3D scene contexts, including egocentric video clips and bird-eye view images, have also been deployed. Notably, researchers have explored the potential of recent language and vision models, including GPT-3 and Unified QA, to tackle the SQA3D task in a zero-shot setting.
Experimental analyses involving state-of-the-art multi-modal reasoning models, such as ScanQA, were conducted to evaluate their efficacy on the SQA3D task against established benchmarks. The outcomes clearly establish SQA3D as a demanding task that necessitates complex reasoning about the agent’s environment for accurate question answering.
4)L3MVN [59] introduces a pioneering module framework that capitalizes on large language models to enhance visual target navigation. The framework aims to solve the challenge of autonomous exploration in unfamiliar settings, where a robot, provided with an object’s name, must proficiently navigate a 3D space to locate that object.
The architecture comprises two principal modules: a language module and a navigation module. The former handles natural language instructions, generating a semantic map embedded with general physical world knowledge. The latter employs this semantic map to guide robotic exploration, deducing the semantic pertinence of visible frontiers and opting for the most cost-efficient maneuvers.
L3MVN’s cardinal innovation resides in its utilization of large language models, specifically GPT-3, to imbue the system with common-sense reasoning for object location tasks. GPT-3, a leading-edge language model, undergoes fine-tuning on a limited set of task-specific data to produce a semantic map replete with object-searching-relevant information.
The theoretical underpinning of L3MVN is anchored in the notion that human adeptness in intricate and dynamic environments stems partially from inherent comprehension of the physical world. This common-sense knowledge can be transferred to robots via large language models, enhancing their operational efficiency in unknown terrains.
Efficacy evaluations of L3MVN were conducted on two comprehensive photorealistic 3D scene datasets, namely Gibson and HM3D. Experimental outcomes indicate substantial enhancements in both the success rate and efficiency of visual target navigation, while obviating the need for expansive learning processes. Additional ablation studies corroborate the efficacy of our language models and cost-utility exploration in facilitating more efficient navigation.
Figure 3: The second type utilizes LLMs to analyze incoming visual or textual data to extract goal-relevant information, upon which exploration policies subsequently generate appropriate actions to guide agent movement. LLMs, by acquiring information from text and other visual models processed through images, perform semantic understanding rather than planning. They extract key information sequences such as targets and locations, and hand them over to exploration algorithms. The exploration algorithms generate actions to guide agents in navigation to complete tasks.
5)ESC [47] (Exploration with Soft Commonsense Constrai-nts for Zero-shot Object Navigation) presents a groundbreaking approach to zero-shot object navigation by leveraging commonsense knowledge from pre-trained models for open-world navigation. Zhou et al. assert that effective object navigation hinges on two essential faculties: (1) Semantic scene understanding, crucial for recognizing objects and rooms, and (2) Commonsense reasoning, needed for logical inferences about probable locations of target objects based on general knowledge.
Previous zero-shot object navigation methods often fall short in commonsense reasoning. They either necessitate training on distinct goal-oriented navigation tasks and scenarios or depend on basic exploration heuristics. These approaches underperform when encountering novel objects or settings. Unlike these methods, ESC employs pre-trained visual and language models for open-world, hint-based grounding via the GLIP model [60], enabling scene comprehension and object grounding. Large-scale image-text pre-training further facilitates generalization to new objects.
Moreover, ESC deploys a pre-trained commonsense reasoning language model to deduce interrelationships between rooms and objects, applying this contextual data for spatial and object reasoning. However, translating this inferred commonsense knowledge into practical actions remains a challenge. Given that relationships between entities tend to be probabilistic rather than absolute, ESC utilizes Probabilistic Soft Logic (PSL [61]) to formulate ”soft” commonsense constraints. These are amalgamated with conventional exploration techniques like Frontier-Based Exploration (FBE) for informed zero-shot decisions on subsequent exploration frontiers.
By setting a new technical work and pioneering the use of pre-trained commonsense knowledge in LLMs for object navigation, ESC paves the way for future research. This could include extracting intricate commonsense relationships, such as spatial connections between rooms, from LLMs. Additionally, future work might explore relaxing the zero-shot stipulations to permit limited fine-tuning in the learning of frontier selection strategies.
6)NavGPT [54] is an instruction-following visual navigation agent founded on Language Learning Models (LLMs). It aims to explore the reasoning capabilities of GPT models in intricate, embodied contexts via zero-shot sequential action prediction. Zhou et al. conducted extensive experiments to show that NavGPT excels in advanced navigation planning, including the dissection of instructions into sub-goals, the incorporation of common-sense knowledge pertinent to navigation tasks, landmark identification in observed scenes, navigation progress tracking, and adjustments to plans based on unexpected developments. The study also revealed LLMs’ aptitude for generating precise navigation instructions from observations and actions, as well as mapping accurate top-down metric trajectories from navigation history.Despite its robust capabilities, NavGPT’s performance in zero-shot Room-to-Room (R2R) tasks does not yet rival that of specialized models. To address this, they advocate the use of LLMs with multi-modal inputs as visual navigation agents and suggest applying LLMs’ explicit reasoning to learning-based models to boost performance.
NavGPT was introduced as a fully automated, LLM-based system designed for language-guided visual navigation. The system accommodates multi-modal inputs, unrestricted language guidance, interactions with open-world settings, and maintains a navigation history. Utilizing GPT-4 for intricate planning, NavGPT demonstrates proficiency in sequential action prediction. It excels in subdividing instructions into sub-goals, integrating navigation-relevant common-sense knowledge, recognizing landmarks in observed settings, tracking ongoing navigation progress, and making plan adjustments based on unexpected events. Furthermore, NavGPT demonstrates spatial and historical awareness by generating trajectory instructions and plotting navigated viewpoints in an overhead view.
Their exhaustive experimentation revealed LLMs’ remarkable abilities for intricate navigation planning, including subdividing instructions into various sub-goals, incorporating relevant common-sense knowledge, recognizing landmarks in observed settings, continually monitoring navigation progress, and adjusting plans based on unforeseen occurrences. The study also confirmed that LLMs can construct navigation trajectories on metric maps and regenerate navigation instructions, thus demonstrating historical and spatial awareness in navigation assignments.
A significant bottleneck in NavGPT is the information loss incurred when translating visual signals into natural language and condensing observations into historical records. Consequently, Zhou et al. propose future research directions that include employing LLMs with multi-modal inputs and navigation systems that leverage the advanced planning, historical, and spatial awareness capabilities of LLMs.
Figure 4: This figure is an example diagram for Planning.
7)VELMA [62] is an LLM-based embodied intelligence agent designed for urban VLN in Street View settings. The agent navigates based on human-generated instructions, which include landmark references and directional cues. Schumann et al. developed VELMA to use verbal cues, contextualized by visual observations and trajectories, to determine subsequent actions. The system identifies landmarks from human-authored navigation instructions and employs CLIP to assess their visibility in the current panoramic view, achieving a linguistic representation of visual information.
LLMs have been increasingly utilized as reasoning engines for embodied agents in diverse applications, such as domestic robotics, video games, and indoor navigation. However, the challenge of interfacing optimally with interactive visual environments persists. This study addresses the issue through verbalization embodiment. Here, the agent’s task, trajectory, and visual observations are verbalized in natural language, allowing LLMs to comprehend navigation instructions and execute corresponding actions.
The paper introduces a landmark scorer to assess landmark visibility within panoramic images. This scorer calculates similarity metrics between textual descriptions of landmarks and their visual representations, using CLIP models. Each landmark receives a normalized score based on visual similarity. If this score exceeds a predefined threshold, the landmark is deemed visible.
The landmark classification is unsupervised, given the absence of ground-truth labels for visibility. The scorer’s threshold is its only tunable parameter. The agent also assesses views to its current left and right orientations. Each directional view is linked to specific textual information. Visible landmarks, along with their associated directional texts, are fed into a language expresser, which then generates verbal descriptions of the environment.
These verbal observations are of two kinds: street intersections and landmarks visible within the current view. These strings encode information about the number of outgoing edges at the current node, the names of visible landmarks, and their associated directional texts.
The study evaluated the performance of the proposed LLM agent in both fine-tuning and contextual learning scenarios. A GPT-3-based landmark extractor was used for a single run and its extracted landmarks were employed in all subsequent experiments. For the landmark scorer, the CLIP-ViT-bigG-14-laion2B-39B-b160k model was chosen as the CLIP model.
Figure 5: This figure is an example diagram for semantic understanding.
8) $A^{2}$av [63] is an action-aware vision-and-language navigation method that consists of an instruction parser and an action-aware navigation strategy. The instruction parser, which employs large language models such as GPT-3, decomposes complex navigation instructions into a sequence of object-specific navigation sub-tasks that relate to particular actions. These sub-tasks necessitate that the agent identify objects and navigate to target locations based on associated action requirements. Despite advancements in VLN, current approaches have failed to sufficiently account for diverse action requirements within instructions, such as ”continue beyond” and ”exit”, thus ignoring the agent-object relationship in the scene.
To tackle this problem, the agent must accurately identify and execute the action requirements associated with each milestone. However, the complexity of grammatical structures and the diversity of action expressions in instructions make understanding these requirements challenging. Additionally, the development of navigation strategies capable of meeting these action requirements without annotated path instructions remains an open question.
$A^{2}$av employs the reasoning capabilities of large language models to parse instructions into action-specific object navigation sub-tasks. In these sub-tasks, the agent identifies landmarks and navigates based on pertinent action requirements. Addressing the limitations of existing zero-shot navigation agents, Chen et al. identified five fundamental action requirements and developed distinct navigators for each. This approach, which obviates the need for manual path instruction annotations, yields more accurate and interpretable navigation outcomes. The system utilizes GPT-3 to parse all landmarks and their associated action requirements, effectively decomposing the instruction into a series of sub-tasks specific to actions.
To execute these sub-tasks sequentially, $A^{2}$av adopts a zero-shot learning approach, comprising five action-specific navigators. While parsing instructions, GPT-3 predicts the ”ACTION”, which typically aligns with one of the predefined ”ACTION” for a sub-task, allowing direct mapping to the corresponding sub-task type. Chen et al. encode all predicted and predefined ”ACTION” using BERT [64] and select the sub-task with the highest cosine similarity score as the predicted sub-task.
For action-aware navigation, Chen et al. fine-tuned ZSON [65] to individually train five distinct action-specific navigators. In each specific action dataset, the positions of landmarks and paths reveal features pertinent to the respective action requirement. In any given task, the current sub-task type is first identified, and the corresponding action-specific navigator is selected to predict low-level actions. The process iterates until either all sub-tasks are completed or the maximum step limit for the episode is reached.
In summary, $A^{2}$av effectively addresses action requirements within VLN tasks by breaking them down into action-specific object navigation sub-tasks. It achieves superior zero-shot VLN performance on the RxR and R2R datasets and outperforms state-of-the-art supervised methods on RxR. Additionally, $A^{2}$av excels in navigating according to instructions with specific action requirements, underscoring its potential in scenarios requiring human-agent interactions.
Table 2: Performance Comparison (Part 1)
Benchmark Habitat RoboTHOR Gibson
SPL SR SPL SR SPL SR
CoW 6.3% - 10% 16.3% - -
ZSON - - - - 28% 36.9%
L3MVN - - - - 37.7% 76.1%
VLMAP 6% 2.5% - - - -
ESC - - 22.2% 38.1% - -
9)MiC [55] (March in Chat) is an environment-aware instruction planner that employs an LLM for dynamic dialogues, specifically designed for the REVERIE dataset. The architecture of MiC is trifurcated into three modules: Generalized Object-and-Scene-Oriented Dynamic Planning (GOSP), Step-by-step Object-and-Scene-Oriented Dynamic Planning (SODP), and Ro-om-and-Object Aware Scene Perceiver (ROASP).
Initiating with GOSP, it queries the LLM to ascertain the target object and its probable locations, subsequently generating a rudimentary task plan. The prompt for SODP is tripartite: the first part utilizes ROASP for scene perception, acquiring room types and visible objects, and translates this information into a natural language description. The second part involves the generation of fine-grained step-by-step instructions based on the selected strategy. The final part includes previously generated instructions. These elements are concatenated and input into the LLM, which then produces detailed planning instructions for the ensuing step.
Concerning ROASP, it not only classifies the room type but also identifies the visible objects in the agent’s immediate environment. Rather than employing separate classifiers and detectors for predicting room types and visible objects, ROASP leverages the CLIP model. With CLIP’s robust zero-shot image classification capabilities, ROASP efficiently performs both tasks. Specifically, it extracts room-type labels from MatterPort3D and object-type labels from REVERIE, then utilizes CLIP for feature extraction.
In operational terms, during a task, GOSP identifies the target object by querying the LLM and utilizes the LLM’s extensive world knowledge to deduce the object’s probable location. Subsequently, ROASP interprets the current scene and prompts the LLM to generate detailed step-by-step plans for the next navigation action. If ROASP discerns a room change, SODP re-queries the LLM to generate a fresh set of fine-grained instructions, aligning them with prior LLM responses. The agent then continues, following these dynamically generated instructions, executing GOSP once and iterating through ROASP and SODP until task completion.
In summary, MiC provides concise high-level instructions to VLN agents, enabling REVERIE agents to dynamically interact with the LLM to formulate plans for upcoming steps. This dynamic planning methodology posits a promising avenue for future research and development.
10)SayNav [53] is a groundbreaking framework designed to harness the common-sense knowledge encapsulated in Large Language Models (LLMs) for generalizing intricate navigation tasks in expansive and unfamiliar environments. SayNav employs a unique grounding mechanism that incrementally constructs a 3D scene graph of the explored territory, subsequently feeding this information to LLMs. This results in the generation of high-level navigation plans that are context-aware and practically implementable, executed thereafter by a pretrained low-level planner as a series of short-distance point-goal navigation sub-tasks.
Incorporating a two-tiered planning architecture and relying on the ProcTHOR [66] dataset, SayNav’s high-level planner extracts subgraphs from the comprehensive 3D scene graph, concentrating on spatial relations in the immediate vicinity of the agent. These subgraphs are converted into textual prompts and presented to the LLM, which in turn produces short-term, stepwise instructions in pseudo-code format. The plans thus generated recommend efficient search strategies, ranking various locations in the room based on their likelihood of containing target objects.
The low-level planner is tasked with generating brief motion control commands. It operates on RGBD images (320x240 resolution) and the agent’s pose data, outputting basic navigational commands in alignment with standard PointNav configurations. To unify the operations of the high-level and low-level planners, SayNav treats each pseudo-code instruction from the LLM as a short-distance point-goal navigation sub-task. The 3D coordinates of the objects specified in each planned step determine the target points for these sub-tasks.
SayNav distinguishes itself by its ability to dynamically generate navigation instructions and iteratively refine its future steps based on newly collected data. Performance metrics on multi-object navigation tasks demonstrate its substantial advantage over Oracle-based PointNav methods, confirming its proficiency in dynamically planning and executing object-finding tasks in large, unfamiliar environments. Moreover, SayNav has proven its efficacy in generalizing from simulated to real-world conditions. In conclusion, SayNav emphasizes the utilization of human-like reasoning and advanced semantic understanding for achieving efficient and adaptive navigation, especially in complex or uncharted settings.
Table 3: Performance Comparison (Part 2)
Benchmark HM3D MP3D REVERIE R2R
SPL SR SPL SR SPL SR SPL SR
ZSON 12.6% 25.5% 4.8% 15.3% - - - -
CLIP-Nav - - - - 4.56% 15.2% - -
L3MVN - - 23.1% 50.4% - - - -
OVRL 29% 64% - - - - - -
ESC 22% 38.5% 13.7% 28.6% - - - -
NavGPT - - - - - - 29% 34%
A2Nav - - - - - - 11% 22.6%
MIC - - - - 41.97% 55.74% - -
SayNav 34% 60.32% - - - - - -
LLMs have demonstrated significant contributions to navigation tasks, primarily through their enhanced capabilities in environmental perception and decision-making. These models integrate advanced language and image processing technologies to provide sophisticated environmental perception and decision support, crucial for tasks requiring a deep understanding of the environment and quick, accurate decision-making. Unlike traditional navigation models that are limited to predefined rules and instructions and offer limited interaction modes, LLMs excel by enabling more dynamic and flexible route planning. Researchers are continuously exploring various applications of LLMs in embodied navigation systems to further enhance their effectiveness in navigation tasks. These applications extend beyond traditional navigation functionalities to include augmented reality perception, multimodal interaction, procedural generation, and automated decision support.
LLMs significantly enhance the environmental perception capabilities within embodied navigation systems through their robust language and image processing abilities. This includes the recognition and interpretation of environmental features and their utilization to make optimal decisions. For example, LLMs can improve navigation performance by analyzing visual information or better understand ground information through the analysis of surface processes and interactions. In terms of decision-making, LLMs are adept at adjusting travel direction and speed based on current location and target information, enabling efficient navigation in unknown or complex environments. They adjust their route based on real-time data and target information to adapt to environmental changes and effectively complete tasks [67].
Given the rapid development of LLMs, several studies have explored methods and practices to enhance traditional navigation tasks using these technologies. LLMs can serve as decision-making components in complex scenarios within autonomous driving systems, by designing cognitive paths and developing algorithms to translate LLM decisions into executable driving commands. This approach not only improves performance in single-vehicle tasks but also handles complex driving behaviors and multi-vehicle coordination, thanks to the common-sense reasoning capabilities of LLMs [68].
For instance, the NaviLLM model [69] supports the execution of multiple navigation tasks such as visual, auditory, and language navigation through pattern-based instructions, enabling a unified model to perform diversely. Some studies have proposed integrating LLMs with visual encoders to enhance the model’s intuitive understanding of the open world, allowing the model to process richer visual information and generate multimodal feedback. This integrated approach not only enhances the model’s interactive capabilities but also improves its performance in open-world scenarios [70]. The powerful reasoning capabilities of LLMs significantly enhance the user experience in embodied navigation, for instance, by utilizing LLMs to interpret user commands in complex or emergency situations, ensuring that the autonomous driving system aligns with user intentions [71]. Simultaneously, by optimizing spatial perception abilities, the system provides more precise spatial location information or scene diagrams, allowing for a more accurate understanding of and response to the user’s spatial position and objectives. Additionally, the DynaCon system [72] utilizes LLMs for context-awareness and dynamic adaptability, enhancing the embodied navigation system’s capabilities in situations requiring a deep understanding of the environment and quick, accurate decision-making. Another study [73] explored the capability of LLMs to generate control scripts to support automated learning of new entries and use natural language supervision to predict the matching of labels with images, facilitating the zero-shot transfer of models to downstream tasks, thereby offering conveniences for the execution of navigation tasks.
3.3.2 Other Models
In addition to the LLM-based navigation models discussed in the previous section, this section compiles several outstanding non-LLM navigation models. These models have also achieved excellent performance through their unique approaches.
1)CoW [74] (CLIP on Wheels) by Gadre et al. [74] is an innovative approach that adapts zero-shot visual models like CLIP to embodied AI tasks, particularly object navigation. The framework involves an agent identifying a target object in unfamiliar environments, defined through text. The key concept is dividing the task into zero-shot object localization and exploration.
However, applying CLIP directly presents challenges. Firstly, CLIP struggles with precise spatial localization crucial for steering the agent. Secondly, CLIP, being static image-based, lacks mechanisms for guiding exploration. Finally, conventional fine-tuning of CLIP might reduce its robustness and generalizability.
To address these, the CoW framework avoids fine-tuning and uses three techniques for object localization: Gradient-based (using CLIP gradients for saliency mapping), k-Patch-based (discretizing the image into sub-patches for individual CLIP model inference), and k-Language-based (matching the entire image with various captions for location information). For exploration, CoW leverages depth maps with two methods: Learning-based (incorporating a GRU, linear actor, and critic heads) and Frontier-based (a top-down map expansion approach) [75].
CoW effectively repurposes zero-shot image classification models for embodied AI, achieving notable success rates by selecting appropriate methods for each task component. While user-specified target exploration is promising, real-world application remains the ultimate test.
2)ZSON [65] introduces an innovative methodology for instructing virtual robots to navigate unfamiliar terrains and identify objects without pre-existing rewards or demonstrations. Distinct from conventional ObjectNav techniques, ZSON exploits image-goal navigation (ImageNav) to transcode goal images into a multimodal, semantic embedding space. This allows for the scalable training of semantic-goal navigation (SemanticNav) agents in unannotated 3D settings. The underpinning theory of ZSON hinges on the principle of semantic similarity. By encoding goal images as semantic embeddings, the approach enables agents to navigate toward objects predicated on their semantic likeness to the goal image. This strategy assumes that objects bearing semantic resemblance to the goal image are likely located in similar spatial contexts.
In the implementation phase, the first step entails using CLIP for pre-training to produce semantic embeddings of image targets. These embeddings encapsulate intricate semantic details about the objectives. Subsequently, a SemanticNav agent undergoes training through reinforcement learning. The agent ingests egocentric RGB observations and semantic target embeddings as input variables and utilizes a ResNet-50 encoder along with a policy network to forecast actions.In performance evaluation, the team executed extensive experiments on three ObjectNav datasets: Gibson, MP3D, and HM3D. Compared to previous zero-shot works, their method has shown a significant improvement.
To conclude, the approach put forth by the researchers yielded commendable success metrics across the three ObjectNav datasets, constituting a significant divergence from traditional ObjectNav strategies. The model is scalable, zero-shot, and well-suited for application in unannotated 3D worlds, establishing it as a viable avenue for open-world ObjectNav implementations.
3)VLMAP [76] presents an innovative spatial map representation that integrates pretrained visual-language features with a 3D environmental reconstruction, facilitating natural language map indexing without the need for extra labeled data. The central innovation is its capacity to intrinsically combine visual-language attributes with the 3D environmental blueprint, empowering robots to execute spatially precise navigation via natural language directives.
The methodology for VLMAP consists of creating a spatial visual-language map capable of direct landmark or spatial reference localization via natural language. Employing readily available visual-language models and conventional 3D reconstruction libraries, VLMAP is assembled by merging visual-language characteristics with a 3D environmental model, thereby allowing natural language map indexing without supplemental labeled data.
VLMAP advances the fields of computer vision, natural language processing, and robotics by unifying visual-language features with 3D reconstructions for precise, natural language-guided robotic navigation. The incorporation of visual-language features leverages recent progress in deep learning, while the 3D reconstruction component builds on established principles of 3D point cloud processing. The natural language capabilities are supported by an expanding corpus of research in natural language processing and human-robot interaction.
Experimental evaluations of VLMAP substantiate its efficacy in enabling robots to navigate spatially precise paths through natural language commands. Tests in a simulated environment affirm the system’s capability to convert natural language directives into a sequence of open-vocabulary navigation objectives that can be directly pinpointed on the map. Further, the adaptability of VLMaps for use among robots with disparate configurations to generate real-time obstacle maps is also demonstrated.
Table 4: Comparison of Works This table provides a comparison of different Works, the design structure, and their respective applications. It’s worth noting that each wok uses different evaluation criteria, which are not listed here. For specific details on each work, please refer to the citations at the end of the article.
Type Works Multimodal Design Structure Specific Applications
LLM-based Models LM-Nav [56] ViNG,CLIP,GPT-3 Translate user commands into a list of landmarks
SQA3D [58] CLIP,BERT,GPT-3 Contextual understanding, zero-sample model exploration
L3MVN [59] RoBERTa Provide common sense for object searching
NavGPT [54] GPT-4 Integrate common sense knowledge and generate actions
VELMA [62] CLIP,GPT Language understanding and generate actions
CLIP-Nav [57] CLIP,GPT-3 Instruction decompose
MiC [55] GPT Planning and generate actions
ESC [47] Deberta Soft common sense constraints and cene understanding
$A^{2}$av [63] GPT-3 instruction decompose
SayNav [53] GPT-3.5-turbo,GPT-4 Dynamically generate step-by-step instructions
Other Models CoW [74] CLIP Object localization,Exploring the environment
ZSON [65] CLIP,ResNet50 Generate semantic image target and
VLMAP [76] LSeg,CLIP Visual language pixel level embedding and 3D map location fusion
OVRL [77] ViT,LSTM Process RGB observations and predict actions
4)OVRL [77] introduces a neural network architecture specifically aimed at visual navigation challenges. Notably, the architecture is composed of task-agnostic elements such as ViTs, convolutions, and LSTMs, obviating the need for task-specific modules. The standout feature of OVRL-V2 is a novel compression layer designed to preserve spatial details in visual navigation assignments. This layer compresses high-dimensional image attributes from the ViT model into low-dimensional features, maintaining the essential spatial context.
The incorporation of the compression layer is grounded in the theoretical premise that retaining spatial information is vital for navigation endeavors, and this layer effectively conserves this attribute. The model further employs self-supervised learning for pretraining, which enhances its generalization performance across new environments.
Architecturally, the paper outlines a universal agent structure consisting of a visual encoder, a goal encoder, and a recurrent policy network. The visual encoder utilizes a ViT-based module for RGB data processing. This output is then amalgamated with a goal representation and channeled through a recurrent LSTM policy network to predict actions. Training methods vary between tasks: IMAGENAV agents are trained via reinforcement learning and DD-PPO [78], while OBJECTNAV agents employ human demonstrations and behavior cloning for training. The visual encoder can either be trained de novo in an end-to-end fashion or pretrained via the MAE algorithm.
Experimental validations affirm OVRL-V2’s state-of-the-art performance in both IMAGENAV and OBJECTNAV tasks, registering success rates of 82.0% and 64.0%, respectively. These results signify marked advancements over preceding methodologies, corroborating the efficacy of both the proposed architecture and its associated training techniques.
Traditional rule-based navigation algorithm (such as Theta* [79], Lifelong Planning A* [80], Multi-Agent Path Finding (MAPF) [81])s and contemporary navigation algorithms each have their strengths and weaknesses. Traditional algorithms have clear rules and stable performance, making them suitable for simple and common scenarios. They have low computational complexity, making them ideal for resource-limited devices and environments. However, the design of these algorithms’ rules cannot cover all possible scenarios, leading to poor flexibility and scalability. Additionally, traditional algorithms struggle to integrate multiple data sources, limiting their ability to handle complex scenarios.
In contrast, modern navigation algorithms, although not entirely based on large language models (LLMs), achieve multimodal data integration through the application of other visual language models. For example, by using weak rules (such as boundary exploration), these algorithms demonstrate strong adaptability, offering new ideas for non-LLM navigation models and proving effective in practical applications.
LLM-based navigation algorithms have the capability to quickly process large amounts of real-time data and support natural language interaction, significantly improving navigation efficiency. Their adaptive and self-learning capabilities surpass those of traditional algorithms. However, LLM algorithms require substantial computational resources and rely on high-quality, extensive data. When handling user data, it is essential to ensure privacy and security to mitigate potential risks.
In summary, traditional rule-based navigation algorithms still have advantages in specific scenarios and resource-constrained environments. However, modern navigation algorithms exhibit greater adaptability and the ability to handle complex scenarios by integrating multimodal data and applying weak rules. LLM-based navigation algorithms offer significant advantages in adaptability, natural language processing, and real-time data processing, but they require careful consideration of computational resources and data security issues.
3.3.3 Comparison
Evaluation metrics play a crucial role in assessing different models, serving as key indicators of model performance. The evaluation metrics for navigation tasks not only focus on Success Rate (SR) and Path Length (PL) but also require corresponding measures to evaluate the consistency between the navigation path and the instructions. Success-weighted Path Length (SPL) is a metric used to evaluate the quality of paths in navigation and path planning tasks. It combines the success rate of the path with the relationship between the actual path length and the ideal path length, thereby providing a comprehensive assessment of path quality. SPL reflects the efficiency and effectiveness of path planning. It considers the comparison between the actual path length and the ideal path length in navigation tasks while weighting the success rate of the path. SPL is widely used in fields such as robotic navigation, autonomous driving, and other areas involving path planning. The calculation of SPL is as follows:
$\text{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{L_{i}}{\max(P_{i},L_{i})}$
In the formula, N is the total number of navigation tasks, $S_{i}$ is the success indicator where $S_{i}$ = 1 if the task is successful and $S_{i}$ = 0 otherwise, $L_{i}$ is the shortest path length from the start point to the target point for the task, and $P_{i}$ is the actual path length for the task.
We have compared the works mentioned above, evaluating aspects such as the large models employed, application domains, and additional features. These comparisons are summarized in Table 4.At the same time, we have divided the aforementioned works into two categories: Planning, as illustrated in fig.4, and Semantic Understanding, as shown in fig.5. Corresponding examples are provided for each category. Furthermore, we have analyzed the performance of these works across various datasets, the details of which are provided in Tables 2 and 3. It is important to clarify that comprehensive evaluation was unfeasible for some works due to the unavailability of public code. Consequently, the results are based solely on performance metrics reported in the corresponding publications. For an in-depth understanding of each work, readers are directed to the cited literature.
Table 5: Evaluation Metrics The metrics of vision-and-language navigation [7].
Evaluation criteria Definition Formula
Path length Length of navigation trajectory from start to stop position $\sum_{\mathbf{v}_{i}\in V}d(\mathbf{v}_{i},\mathbf{v}_{i+1})$
Navigation error Distance between predicted path endpoint and reference path endpoint $d(v_{t},v_{e})$
Ideal success rate The probability that any node in the predicted path is within a threshold distance from the endpoint of the reference path $\mathbb{I}\left[\left(\min_{\mathbf{v}_{i}\in V}d(\mathbf{v}_{i},v_{e})\right)\leq d_{th}\right]$
Navigation success rate The probability that the stopping position is within 3 meters of the end point of the reference path $\mathbb{I}\left[NE(v_{t},v_{e})\leq d_{th}\right]$
Success rate based on path weighting Navigation success rate based on path length weighting $SR(v_{t},v_{e})\cdot\frac{1}{d_{gt}}\max\left\{PL(V),d_{gt}\right\}$
Weighted coverage score Coverage and length scores of predicted paths relative to reference paths $PC(P,R)\cdot LS(P,R)$
Based on dynamic time warping weighted success rate Success rate of predicted path to reference path based on dynamic time warping and spatial similarity $SR(v_{t},v_{e})\cdot\exp\left(\frac{1}{|R|}\min_{w\in W}\sum_{(i_{k},j_{k})\in w}d(r_{i_{k}},q_{j_{k}})\right)$
3.4 Other Applications of LLMs in Embodied Intelligence
1)VoxPoser [3] focuses not on embodied navigation works but on embodied AI. VoxPoser is a novel approach that utilizes LLMs to guide robots in performing various tasks without the need for predefined motion primitives. Through LLMs, it can extract crucial cues and constraints from natural language instructions to generate robot motion trajectories. For example, when given the command ”Open the topmost drawer, but be careful of the vase,” LLMs can understand the need to grasp the drawer handle and pull it outward while avoiding the vase. Subsequently, by assembling 3D value maps, robots can execute these actions in the real environment without the need for additional training.
Unlike traditional methods, VoxPoser not only addresses the issue of instruction execution but also generates robust robot trajectories in dynamic environments. It leverages the code-generation capabilities of LLMs to translate language instructions into Python code, which invokes perception APIs to obtain spatial information about objects in the environment, thus creating 3D voxel maps. These maps serve as the objective function for motion planners, guiding robots in action execution. The uniqueness of this approach lies in its ability to allow robots to perform tasks based on natural language instructions in unknown scenarios, thereby achieving zero-shot generalization to open instructions.
Overall, VoxPoser provides a new paradigm for robots to perform complex tasks, making full use of the capabilities of language models to interpret and execute natural language instructions. Its application potential extends beyond robot operation to other domains such as autonomous driving vehicles or smart home systems.
2)ALFRED [82] (Action Learning From Realistic Environments and Directives) serves as both a benchmark and dataset for the training and evaluation of models across a spectrum of embodied intelligence tasks. The central focus of this benchmark is to examine the efficacy of AI agents in carrying out domestic tasks, based on detailed natural language directives. The dataset comprises 25,743 English instructions that correspond to 8,055 expert demonstrations. Each demonstration encompasses approximately 50 steps, culminating in a total of 428,322 paired image-action sequences.
ALFRED tasks are highly complex, often necessitating that agents execute a chain of sub-tasks involving irreversible state alterations, such as breaking an object or transferring food between containers. Included in the dataset are high-level objectives like ”rinse a mug and place it in the coffee maker”, as well as more granular instructions such as ”walk to the coffee maker on the right”. Robots interacting with human habitats must learn to associate natural language with their surroundings. ALFRED outperforms existing vision and language task datasets by offering a wealth of expert demonstrations in simulation environments that are more intricate in terms of sequence length, action space, and language complexity.
Contrary to the advancements made possible by datasets like MP3D, these resources often fail to address the complexities arising from task-oriented object interactions. ALFRED mandates that agents extract essential information from natural language directives, encompassing both the goals and potential execution steps for tasks. Agents are also required to deconstruct high-level tasks into executable actions or sub-tasks and plan an optimal path for task completion. Additionally, they must identify various objects and environmental features from a first-person perspective and engage in complex interactions with environmental objects, such as picking up items, manipulating switches, or navigating to specific locations.
By offering a visually and physically realistic simulated environment that enables translation from language to action sequences and interactions, ALFRED tackles numerous challenges evident in real-world scenarios. This includes translating human language into robotic actions for household tasks. Models capable of overcoming these challenges will begin to narrow the divide between real-world, language-driven robots and their simulated analogs.
3)PaLM-E [83] (Pre-training across Language and Modality Embodied) serves as a multimodal language model proficient in integrating visual, linguistic, and sensor modalities to facilitate more grounded reasoning for complex real-world challenges. Specializing in robotic planning, visual question-answering, image description, and linguistic tasks, PaLM-E exhibits robust transfer learning capabilities. It trains as a generalized, efficient embodied reasoner by merging embeddings of real-world entities in its dataset, making it skillful in multi-agent decision-making and transfer learning. Experiments indicate that concurrent training on diverse robotic tasks and datasets bolsters its performance and generalization capacity. Moreover, PaLM-E excels in language-conditioned policies and self-supervised learning, enabling both zero-shot and one-shot generalizations.
While LLMs display substantial reasoning capabilities across an array of domains, they are constrained by a grounding issue. Previous research, such as work by Ahn et al. [84], combined the outputs of LLMs with learned robotic policies and affordance functions. However, these efforts were limited as they supplied only textual inputs to the LLM, insufficient for tasks demanding geometric scene understanding. Training LLMs on expansive textual data may yield representations pertaining to the physical world, yet integrating these with real-world visual and sensor modalities is essential for addressing broader, grounded challenges in computer vision and robotics.
PaLM-E augments this by directly assimilating continuous sensor modalities within the agent, thereby enriching the language model’s capacity for informed, sequential decision-making in real-world contexts. Inputs like images and state estimates are encoded in the same latent space as linguistic tokens, and are processed through the self-attention layers of a Transformer-based LLM analogously to text. Driess et al. initiated this process with a pre-trained LLM, incorporating continuous inputs via encoders. These encoders undergo end-to-end training to produce sequential decisions in natural language, interpretable by agents through fine-tuned, low-level policies or by providing specific query responses.
Utilizing the Transformer architecture and drawing from a diverse set of training data that encompasses text, images, and audio, PaLM-E can comprehend and manage multiple information types. It boasts zero-shot learning capabilities, enabling it to tackle previously unencountered tasks via inference and generation. The model accommodates a broad array of inputs like text, images, and audio, rendering it versatile for complex tasks. Its generative prowess extends to the production of natural language text, image descriptions, and beyond, thereby making it applicable across various natural language generation domains.
4)RT-2 [85] (Robotics Transformer 2) serves as an avant-garde model crafted to adapt web-sourced visual and linguistic insights into robotic control functions. Composed of dual core elements—a pre-trained vision-language model and a reinforcement learning-based robotic control policy—RT-2 leverages expansive web data to cultivate comprehensive visual and linguistic awareness. Exhibiting remarkable proficiency in both simulated and authentic environments, the model excels in executing intricate operations such as object manipulation and door-opening tasks.
Contemporary researchers like Brohan et al. contend that LLMs function primarily at a macroscopic level, essentially acting as state machines that dictate orders to independent, microscopic controllers. These low-level controllers often lack access to the rich semantic knowledge harbored within web-scale models during their developmental stages. They probe the feasibility of incorporating pre-trained, large-scale Visual Language Models directly into these microscopic controllers to amplify their generalization and emergent semantic reasoning faculties. In response, they develop visual-language models tailored for open-vocabulary visual question-answering and dialogue, programmed to generate low-level robotic actions, while concurrently addressing broader web-scale visual-language challenges.
Through the refinement of existing models such as PaLM-E [83] and PaLI-X [86] into RT-2-PaLM-E and RT-2-PaLI-X, Brohan et al. engineer high-efficacy robotic schemes that inherit the advanced generalization and emergency response features facilitated by web-scale visual-language pre-training. This unambiguous yet multifaceted strategy elucidates the immense potential for robotic technologies to gain directly from the strides made in visual-language models, thereby aligning the progression of robotics with concurrent advancements in complementary fields.
Nonetheless, deploying large-scale Visual-Language-Action (VLA) models for real-time operations necessitates considerable computational resources, thereby raising issues for scenarios that demand high-frequency control. A promising avenue for ensuing research encompasses the exploration of quantization and distillation techniques, which could facilitate the functioning of these models at elevated speeds or on more economically viable hardware platforms.
3.5 State-of-the-Art Datasets
Recently, with the accelerated advancement of AI technologies, the focus of interest has transitioned from digital interfaces to tangible environments, catalyzing the emergence of a novel research frontier—Embodied AI. As implied by its nomenclature, Embodied AI primarily investigates the dynamic interplay between an autonomous agent and its real-world environment. To execute tasks both successfully and robustly, agents require advanced proficiencies in environmental perception, encompassing, but not limited to, faculties in vision, language, reasoning, and planning. Consequently, the selection of an appropriate dataset for pre-training these agents becomes a critical determinant of their performance.
Current explorations within the domain of Embodied AI encompass a diverse array of applications including navigation, object manipulation, and multimodal learning. Correspondingly, the datasets requisite for these distinct research areas exhibit considerable variation. This section will present an overview of multiple datasets pertinent to embodied intelligence: MP3D [87], TOUCHDOWN [88], R2R [89], CVDN [90], REVERIE [91], RXR [92], SOON [93], ProcTHOR [66], R3ED [94], and X-Embodiment [95]. The chronological development of these datasets is depicted in Figure 6.
Figure 6: The timeline below provides an overview of the key, commonly referenced datasets discussed in this paper, covering a period from 2017 to 2023.
1)MatterPort3D [87] (MP3D) is a comprehensive dataset, predominantly comprising RGB-D scenes. It features 10,800 panoramic views generated from 194,400 RGB-D images and spans 90 expansive architectural settings, primarily focusing on 3D indoor environments. The dataset encompasses various data modalities such as RGB images, depth images, and semantic annotations. Annotations are supplied for facets like scene surfaces, camera poses, and semantic segmentation. Given its voluminous collection of scenes and precise global perspectives, MP3D serves as a pivotal resource for tasks reliant on self-supervision or alternative methods.
As a cornerstone in the field, MP3D has influenced the development of numerous subsequent datasets, which have either adapted or expanded upon its structure. This renders it an invaluable asset for simulating Embodied AI in computational settings. Nevertheless, the dataset’s primary focus on indoor residential environments circumscribes its applicability. Unique to MP3D, each viewpoint incorporates a 360-degree panoramic image, encompassing both depth and color information. The dataset meticulously samples viewpoints at human height throughout these spaces, ensuring camera poses are aligned. Data are amassed across a range of applications, from semantic segmentation to object and regional categorization.
For each object surface, MP3D furnishes multiple views captured from diverse angles and distances, thereby enabling the learning and prediction of surface properties contingent on these perspectives. The dataset places a significant emphasis on visual perception and scene understanding, making it highly versatile for a plethora of computer vision tasks including object detection, scene segmentation, and 3D reconstruction. For Embodied AI research, MP3D stands as an exemplary dataset, offering a data-rich environment that enhances agent-environm-ent interactions.
2)TOUCHDOWN [88] is a specialized dataset designed for navigation and spatial reasoning, utilizing Google Street View to construct an expansive outdoor environment. The dataset encompasses 39,641 panoramic images and 61,319 edges, all originating from New York City. Each viewpoint in the dataset features a 360-degree RGB panorama. To compile this dataset, TOUCHDOWN employs a ”treasure hunt” methodology, involving two roles: the writer and the finder. The writer crafts instructions to enable the finder to navigate toward and locate a specific target, without prior knowledge of the correct route or destination. Upon task completion, the finder documents the followed path and elucidates their navigation choices.
TOUCHDOWN consists of two primary tasks: Navigation and Spatial Description Resolution (SDR). The Navigation task emphasizes spatial reasoning, with instructions focusing on the spatial relationships between the agent and its environment. The SDR task, on the other hand, demands linguistic expertise to decipher relationships among objects and their spatial configurations. Each of these tasks presents its own set of unique challenges.
Contrary to MP3D [87], which is confined to indoor scenarios, TOUCHDOWN predominantly features outdoor settings. This makes it an ideal dataset for training and evaluating models in tasks that require outdoor navigation and spatial understanding.
3)Room-to-Room [89](R2R) is a specialized dataset designed for VLN, extending the data from MP3D to include a substantial corpus of natural language data pertinent to navigation. R2R characterizes a series of navigational paths within these 3D indoor settings, each consisting of a sequence of specific locations and orientations, termed ’viewpoints.’ Accompanying each path are one or more natural language instructions that guide navigation from the origin to the destination.
Unlike MP3D, which is more general-purpose, R2R employs these indoor settings for highly specialized navigation tasks, necessitating a nuanced understanding of natural language and complex decision-making. Navigating intricate indoor terrains often demands capabilities in long-term memory and reasoning. A model utilizing R2R not only needs to comprehend natural language directives but must also correlate them with the visual landscape. Upon this synthesis, the model is required to strategize an efficient route from the starting location to the specified endpoint. Through these multiple dimensions, R2R offers an intricate and demanding research platform for those specializing in vision-and-language navigation.
4)CVDN [90] is an extensive dataset comprising over 2,000 tailored ”human-to-human” dialogues, set in a simulated, photorealistic domestic environment. Developed as an extension of the R2R framework and formulated in English, CVDN aims to cultivate navigation agents capable of functioning in residential and commercial spaces via linguistic directives. Notably, these agents have the ability to pose focused questions during instances of uncertainty regarding subsequent navigational steps.
Furthermore, CVDN can be leveraged to develop agents endowed with specialized environmental knowledge, thereby facilitating automated linguistic guidance for individuals in unfamiliar contexts—such as direction-seeking within an office complex. The high fidelity of the simulated environment ensures that agents trained on CVDN can generalize skills to real-world settings.
Diverging from R2R, dialogues within CVDN typically contain nearly triple the word count and describe paths averaging over thrice the length. Thus, CVDN provides a significant avenue for the exploration of in-situ navigational interactions and holds potential for the training of agents proficient in both navigation and dynamic inquiry within human-centric environments.
5)REVERIE [91] is a comprehensive dataset featuring 10,567 panoramic images spread across 90 buildings and encompassing 4,140 target objects. Accompanying these are 21,702 crowd-sourced instructions, averaging 18 words each. While it shares visual data attributes like RGB images with earlier datasets, REVERIE also incorporates enhanced semantic particulars related to the Goal Object, such as color, along with natural language directives.
Distinct from conventional VLN datasets, REVERIE includes a myriad of complex and diverse indoor environments. Moreover, its instruction set diverges from the step-by-step commands typical of R2R datasets to offer semantic-level directives more congruent with real-world tasks and human-agent interactions. For example, instead of sequential directives like ’Start at the entrance and proceed to the red sofa,’ REVERIE provides semantically rich tasks such as ’Fold the towel in the bathroom with the fishing theme.’
This semantic depth, however, introduces a layer of complexity due to potential ambiguities (e.g., ’Fold the towel in the bathroom with the fishing theme’ could signify either ’Fold the fishing-themed towel’ or ’in the fishing-themed bathroom’). Consequently, agents trained on the REVERIE dataset require enhanced long-term memory, semantic comprehension, and environmental reasoning abilities. Such attributes make REVERIE particularly suited for remote entity referencing and multi-task learning endeavors. In summary, REVERIE offers a multifaceted, challenging platform conducive to advancements in VLN and multimodal learning.
6)RxR [92](Room-Across-Room) is an expansive dataset tailored for Multilingual VLN tasks. As an extension of the Room-to-Room (R2R) dataset, it inherits substantial visual and navigational data, thereby enriching the research landscape with multilingual capabilities. The dataset features natural language navigation instructions in multiple languages, including but not limited to, English, German, and Spanish. This linguistic variety significantly broadens the scope for multilingual VLN endeavors.
Furthermore, RxR introduces dense spatiotemporal grounding, which entails comprehensive action execution and natural language interpretation directives, calibrated to specific temporal intervals and spatial coordinates within the navigation trajectory. Similar to REVERIE, RxR augments the dataset with additional layers of semantic data pertaining to target objects and multiple languages.
This compounded complexity poses new sets of challenges: the navigation agents must master multilingual understanding, spatiotemporal reasoning, and long-term planning. Owing to its linguistic diversity, RxR serves as a multifaceted, demanding research substrate, suitable for advancing the state-of-the-art in multilingual VLN.
7)SOON [93], also known as the FAO dataset, is a language-guided navigation dataset grounded in 3D housing environments featuring authentic image panoramas. The dataset encompasses 90 distinct residential settings, each paired with a natural language instruction that delineates a target location. Unique to SOON, the construction of these instructions is organized around five subtasks: delineation of target attributes, explication of relationships between the target and neighboring objects, exploration of the target’s regional setting, scrutiny and description of adjacent areas, and the subsequent condensation of these elements into a three-sentence narrative.
In addition to offering an extensive array of 3,848 instructions characterized by a diverse vocabulary of 1,649 words, the SOON dataset furnishes high-fidelity annotations, which include 3D environmental models, RGB-D imagery, semantic segmentation, and depth data. Remarkably, the average instruction length in SOON stands at 38.6 words, surpassing the 26.3-word and 18.3-word averages observed in REVERIE and R2R, respectively. Most instructions oscillate between 20 and 60 words, affording robust descriptive potential.
In contrast with other datasets, SOON stands out for its elevated linguistic intricacy and ecological diversity, coupled with a rich, realistic representation of environmental contexts. Consequently, the SOON dataset provides an exhaustive and reliable substrate for advancing research in the field of vision and language-guided navigation.
8)ProcTHOR [66], also known as PROCTHOR-10K, is a dataset comprising 10,000 algorithmically generated 3D residential scenes specifically tailored for the training and assessment of embodied AI agents. The dataset is a byproduct of the PROCTHOR framework, which is designed to facilitate the large-scale generation of diverse, interactive, and customizable virtual environments. Each scene within the dataset is distinct, furnished with a multiplicity of objects, textural variations, and illumination conditions, thereby constituting a rich and complex milieu for embodied agents to operate, interact, and manipulate.
Intended to support an array of embodied AI tasks—ranging from navigation and object manipulation to comprehensive scene understanding—PROCTHOR-10K has been employed for the training and evaluation of cutting-edge embodied AI models. The dataset has demonstrated robust performance across various industry benchmarks, including but not limited to Habitat 2022, AI2-THOR Rearrangement 2022, and RoboTHOR challenges.
9)R3ED [94] is a dataset encompassing over 5,800 point clouds and approximately 22,400 annotated 3D bounding boxes, sourced from seven distinct real-world indoor settings via a dense sampling technique. This dataset aspires to address the limitations inherent in previous embodied AI datasets, which predominantly rely on synthetic data to mimic indoor landscapes and robotic motions. While synthetic data offers convenience, it tends to undermine real-world performance.
Conversely, R3ED offers authentic point cloud data following each agent’s move, furnishing researchers with the opportunity to examine active vision through a more realistic and standardized lens. The point cloud data is meticulously annotated with 3D bounding boxes that categorize 12 types of objects, inclusive of prevalent furniture and appliances. Furthermore, the geometrical relationships among these point clouds are explicitly annotated using a specialized storage path format. To augment the dataset’s utility, the authors have introduced a 3D divergency policy aimed at directing robotic movements for data collection, thereby enhancing visual performance in unfamiliar environments.
Table 6: Comparison of Datasets. This table is a comparison among nine datasets, including seven evaluation methods.
Dataset Size Content Real(Physical)
Panoramic Instruction Point Cloud Human Inclusion
MP3D 10,800 - - View,Annotation
TOUCHDOWN 39,641 - - Images,Edges
R2R - 21,567 - Nav-Instruction
REVERIE 10,567 21,702 - View,Instruction
RxR - 126,000 - Nav-Instruction
SOON - 3,848 - Instruction,Location
ProcTHOR 10,000 - - View
R3ED - - 5800 PointCloud,Bbox
CVDN - - - Dialog
10)X-Embodiment[95] is an embodied robot manipulation dataset, comprises over one million robot trajectories from 22 disparate robot instances, illustrating 527 skills across 160,266 tasks. The dataset amalgamates data from 60 pre-existing robot datasets, sourced from 34 international robotics research labs, and standardizes them into a unified RLDS data format[96]. This format efficiently serializes record files and accommodates a myriad of action spaces and input modalities, from varying numbers of RGB cameras to depth cameras and point clouds, thereby facilitating streamlined data loading across all primary deep learning frameworks. Transformer-based policies trained on this dataset have shown marked positive transfer capabilities across different robotic instances.
Recent developments in machine learning and artificial intelligence indicate that large-scale, diversified datasets are pivotal for the creation of competent AI systems. Such is evident in domains like natural language processing and computer vision, where large, high-capacity models trained on diverse data have effectively addressed a range of downstream applications. This trend has catalyzed the widespread adoption of pretrained models as foundational assets across applications. In the realm of robotics, models have traditionally been specialized for specific applications, robots, and environments. However, there is burgeoning interest in the development of ”generalist” robotic policies capable of efficiently adapting to novel robots, tasks, and settings. Large-scale, multipurpose models, often trained on diverse datasets, consistently outperform task-specific counterparts trained on more limited datasets.
Prompted by the efficacy of pretraining expansive vision or language models on varied data, DeepMind posits that the construction of universally applicable robotic policies necessitates X-embodiment training. This approach utilizes data aggregated from multiple robotic platforms. Although individual robotic learning datasets might be too specialized, their combination provides a more comprehensive representation of environmental and robotic variations. The development of methodologies to optimally leverage this X-embodiment data from diverse labs, robots, and settings is indispensable for the creation of universally applicable robotic policies. While current datasets might not yet match the generalization capabilities exhibited by large language models, the collaborative utilization of such resources has the prospective capacity to achieve comparable coverage. Consequently, DeepMind contends that research into X-embodiment robotic learning is of paramount importance at this stage.
In summary, this section provides an overview of ten seminal datasets, delineating their structures and merits. Table 6 offers a comparative analysis among these datasets, with more exhaustive information available in the References. These datasets provide diverse and rich resources for LLM research in the navigation domain, facilitating the optimization and development of models across various environments and tasks. MP3D offers high-quality 3D indoor environment data, aiding LLMs in understanding complex indoor scenes; TOUCHDOWN provides urban streetscape data suitable for studying natural language navigation in outdoor environments; R2R features real user-generated navigation instructions, enhancing the naturalness and diversity of the data; RXR supports multi-language navigation instructions, improving the model’s cross-language adaptability; REVERIE provides target-oriented navigation task data, promoting fine-grained navigation research; ProcTHOR offers synthetic virtual environments, supporting large-scale model training; R3ED introduces memory modules, supporting long-term navigation tasks; CVDN includes human-to-human dialogues, guiding agents in environmental navigation, suitable for studying human-machine interaction and collaborative navigation. By integrating high-quality data and multimodal inputs, these datasets advance LLMs in natural language processing, visual understanding, and multi-task learning, enhancing the intelligence and practicality of navigation systems. Broadly, both contemporary and forthcoming datasets are gravitating towards the inclusion of real-world 3D data, multi-modal inputs, higher-order instructions, and provisions for human oversight and direction.
4 Challenges and Future Directions
LLMs also face many challenges and limitations in navigation applications. In terms of datasets, each dataset has its unique strengths and weaknesses for different types of navigation and language guidance tasks. Matterport3D, R2R, RXR and REVERIE [87, 89, 92, 91] are suitable for indoor navigation tasks due to their high realism and rich annotations; TOUCHDOWN [88] is suitable for navigating urban environments; ProcTHOR [66] is applicable to the need for controlled and scalable environments; and R3ED [94] is suitable for long-term navigation tasks by introducing a memory module. Taking several datasets as examples, the MP3D dataset comes from real-world indoor environments, covering a wide range of indoor scenes such as houses and offices, with high diversity and containing detailed object labeling and semantic information, which can provide high-quality RGB-D images and 3D reconstruction data, but it also increases the difficulty of model training and inference due to the complexity of the scenes, and requires a high hardware configuration  [87]. R2R and REVERIE extend on the basis of MP3D. R2R’s scene is based on MP3D while providing real user-generated navigation instructions, but the number is limited and not diversified enough; REVERIE not only provides instructions, but also includes descriptions of the target objects, which is suitable for more fine-grained and goal-oriented navigation tasks, but with higher task complexity, and higher requirements for processing and storage. Higher. In contrast to the previous indoor datasets, TOUCHDOWN provides urban street scene data with natural language descriptions to guide navigation tasks, which is suitable for outdoor tasks, but urban scenes are more complex and the annotation quality of natural language descriptions may be inconsistent.ProcTHOR provides synthetic virtual environments, which allow the user to customize the complexity and layout of the environment, and therefore can generate a large number of scenes. CVDN [90] , CVDN contains dialogues between humans to guide the Agent to navigate in the environment, and the task design emphasizes cooperativeness, which is suitable for studying human-computer interaction and cooperative navigation, but the multiple rounds of dialogues and collaborative tasks increase the complexity of model training. the complexity of model training. The existing dataset has a large amount of data, but the diversity about objects is insufficient to meet human needs in specific situations [74].
In terms of modeling, the existing LLM navigation model utilizes the powerful natural language processing capabilities of LLM with extensive knowledge to guide the Agent in various environments. Since LLM is capable of understanding and generating natural language text, it can well simulate human language or accomplish tasks through conversations [55]. In addition, LLMs fine-tuned in various ways can handle multimodal data, such as combining multimodal inputs such as text, images, and speech, which can improve the accuracy of navigation and enhance the user interaction experience. However, LLM requires a large amount of high-quality data to train the model, and insufficient or inaccurate data will affect the navigation effect [62]. Although most of the existing LLM-based navigation models use pre-trained or directly invoked LLMs, they are not directly and task-specific trained, which can also adversely affect the final navigation results. Compared to non-LLM approaches, which do not need to rely on large amounts of natural language data, have relatively low data requirements, rely on image-text matching, and do not have the same scope of knowledge and comprehension as LLMs. LLMs have limitations in their ability to deal with fine-grained navigational instructions and spatial direction inference, and perform especially poorly in tasks that require long-term planning and spatial reasoning. In addition, LLMs are not sufficiently adapted to interactive navigation and show limitations in navigation tasks in environments containing traversable obstacles  [97]. Taken together LLM-based navigation methods are suitable for complex dialog and multimodal tasks, whereas non-LLM navigation methods  [57]perform well in image-text matching and visual tasks, with the exact choice depending on task requirements and resource conditions.
In models like NaVid [98], continuous RGB video frames serve as observational inputs, while instructions function as language inputs. Visual information from each frame is extracted using a frozen observation encoder, while instructions are transformed into language tokens using a text encoder. The information is then segmented into past frame data, current frame data, and navigation information, and Vicuna-7B is used to obtain the next action command. This multimodal fusion approach [99], along with the use of pre-trained models [100], is widely adopted in current LLM-based VLN models. The integration of multimodal information, such as in NaVid, including temporal, visual, and instruction information, can effectively enhance the model’s efficacy, while pre-training methods can alleviate training costs. This is also the primary direction for the future development of LLM-based VLN models.
When exploring the development and integration of LLMs in physically intelligent systems, it is necessary to consider the powerful environment-aware understanding provided by the development of LLMs, and the navigation performance of robots can be significantly improved by integrating LLMs with other modules, such as multi-sensor information fusion techniques. In addition, the predictive capabilities of LLMs can be further enhanced by introducing methods to predict future behaviors, enabling them to better understand and predict future events. For navigation tasks, studies have shown that the navigation ability of robots in complex environments can be effectively enhanced by integrating multimodal inputs such as vision, hearing, and haptics, and by utilizing complex decision-making frameworks [101, 102]. Future research should focus on developing generalized models that can handle multimodal inputs and perform complex tasks, while ensuring that these models remain efficient and flexible across different application scenarios. This includes, but is not limited to, the use of LLMs for multi-task learning and the development of new algorithms to optimize the robot’s behavioral strategies to adapt to changing environments [69]. According to the study, path planning can be performed more efficiently by using A* search algorithms and incorporating human feedback, while targeting in navigation specific behavioral control can be done by using LLMs to generate low-level control commands to directly control the robot’s dynamic motion without task-specific fine-tuning. This can be achieved by collecting cues from the physical environment, thus enabling the robot to perform basic actions such as walking [103]. At the same time, it is important for embodied subjects to form “memories” and to be able to avoid “forgetting” errors. LLMs may miss critical information when executing robot programs. Based on this, cue engineering strategies should be designed to minimize execution errors. By enhancing the multimodal processing capability of LLMs, using multimodal instruction tuning and multimodal learning strategies, the cross-modal dialog capability of LLMs can be greatly improved, and then they can be more capable of navigating tasks. In practice, multimodal command tuning technology can realize more accurate user intent understanding by combining voice commands and image input. By jointly learning data from different modalities such as speech, image and text, a richer and more comprehensive navigation model can be built [104].
In the interaction between embodied intelligence and the actual environment, the issues of data security and user privacy protection are particularly prominent. This is not only a technical issue, but also involves social ethics, legal norms and technological innovation. Meanwhile, the decision-making process of LLMs is often black-box, and it is difficult for developers to fully understand its internal mechanisms. In navigation applications, if users are unable to understand how LLMs make specific navigation recommendations, this may affect their trust and acceptance of the system. In terms of ensuring data security and protecting user privacy, a series of stringent measures must be taken to ensure that user data are not misused, leaked or tampered with. Some methods are used in protecting user privacy, such as training with synthetic data [105] or monitoring models’ outputs [106]. Therefore, privacy protection needs to be taken into account at the design stage, which means that it should be regarded as a fundamental principle rather than an afterthought remedy during the construction of embodied intelligence systems. The security of data collection, transmission, storage and processing should be considered at an early stage of system design, and appropriate encryption, access control and authentication mechanisms should be adopted to ensure the confidentiality and integrity of user data. In addition, a strict data access authority system should be established to limit the scope and purpose of data use to avoid data abuse or illegal use [107].
In the future, LLM-based navigation models may evolve in the following directions:
  1. 1.
    dynamic path optimization, which is no longer bound to static space (e.g. MP3D) but moves towards dynamic space to adjust paths in real time according to dynamic targets in space.
  2. 2.
    optimization of algorithms and model architectures using LLMs trained for a specific task in order to reduce resource consumption and improve efficiency.
  3. 3.
    integration with automated driving technologies. by fusing multimodal data, such as GPS radar, to provide high-precision navigation support while improving the user interaction experience.
On the whole, LLMs have great potential for development in navigation tasks, but still need to solve the challenges of data diversity, fine-grained navigation, spatial reasoning ability and interaction adaptability, and improve their application in complex environments through multimodal fusion and optimization algorithms.
5 Conclusion
This paper delves into the rapidly evolving field of Embodied Intelligence, focusing on the premise that intelligence emanates from the interaction between an agent and its environment rather than being a purely internal, abstract construct. The paper scrutinizes 14 works in embodied navigation, delineating their variances and commonalities. Furthermore, it offers an in-depth analysis of how these works catalyze contemporary advancements in research tasks in embodied AI, shedding light on their underlying design principles and methodologies. Experimental outcomes indicate significant success rate improvements across various work evaluations.
The paper also outlines challenges and limitations inherent to LLMs in the realm of embodied intelligence. These include the absence of a direct linkage between LLMs and the physical world, the necessity for extensive training data, and the complexity of understanding and generating natural language in a diverse array of contexts. Notwithstanding these impediments, several promising research trajectories exist that could propel advancements in embodied intelligence. These encompass the formulation of more robust and adaptable language models, the investigation of innovative training methodologies and architectures, the establishment of standardized works and evaluation metrics, and the critical need for cross-disciplinary collaboration among researchers in AI, robotics, and social sciences.